# Exploring Multidimensional Checkworthiness: Designing AI-assisted Claim Prioritization for Human Fact-checkers

### v2

@@ Abstract.
Given the massive volume of potentially false claims circulating online, claim prioritization is essential in allocating limited human resources available for fact-checking. In this study, we perceive claim prioritization as an information retrieval (IR) task: just as multidimensional IR relevance, with many factors influencing which search results a user deems relevant, checkworthiness is also multi-faceted, subjective, and even personal, with many factors influencing how fact-checkers triage and select which claims to check. Our study investigated both the multidimensional nature of checkworthiness and effective tool support to assist fact-checkers in claim prioritization. Methodologically, we pursued Research through Design combined with mixed-method evaluation.
Specifically, we developed an AI-assisted claim prioritization prototype as a probe to explore how fact-checkers use multidimensional checkworthy factors to prioritize claims, simultaneously probing fact-checker needs and exploring the design space to meet those needs. With 16 professional fact-checkers participating in our study, we uncovered a hierarchical prioritization strategy fact-checkers implicitly use, revealing an underexplored aspect of their workflow, with actionable design recommendations for improving claim triage across multidimensional checkworthiness and tailoring this process with LLM integration.
@@
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Interactive systems and toolsccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing User centered design
@@ 3.1. Frame Research Goals
Figure 1. Our lab-based RtD method integrated into a classic double diamond process (Design Council, 2015). We document our design challenges as the research goals, described in Section 3.1. Steps of Discover and Define, i.e., how we explore and finalize design choices, are documented in Section 3.2. The step of Develop, i.e., how we create and deploy the prototype, is documented in Section 3.3. The step of Deliver, i.e., how we finally evaluate the design to gather use-inspired insights, is documented in Section 3.4. The iterative refinement of RtD evaluation protocol is documented in Appendix E.
@@ 3.3.2. Technical implementation
We implemented our models on the Streamlit server and built the front-end [^8] Streamlit: https://streamlit.io/, AG-Grid: https://www.ag-grid.com/ using Python, AG-Grid JS, and CSS. We used OpenAI’s gpt-3.5-turbo as our LLM to create customized classifiers [^9] While more advanced GPT models now exist, the log probability needed for building the ranking function was only available from gpt-3.5-turbo at the time of our implementation..
@@ 3.4.2. Baseline condition
ID Gender Region Position Years of Experience Organizational Context Language Fact-checked
@@ 4.1.1. Perceived importance
Table 4. Mean (M), standard deviation (SD), and median statistics over participant data for different measures (rows, Section 3.4.4) and checkworthy dimensions (columns). Self-assessment data (top 3 rows) come from 5-point Likert scale answers, while observational data (bottom 3 rows) is drawn from interaction logs for the different UI sliders. Bold results indicate at least two checkworthy dimensions showed statistically significant differences (Friedman test at p ¡ 0.05). We observe significant differences between checkworthy dimensions for “Ease of finding” and “Overall weight” measures. Post-hoc Wilcoxon test results are presented in Appendix H.
@@ 4.3. Targeted versus Abstract Prompting for Defining New Checkworthy Dimensions
I’ve done a lot of reporting on this [VAERS] and I’m familiar with the language, that’s like a huge subsection of COVID-19 claims… prompt like this helps me locate them [quickly].” (P13)
@@ 4.4.1. Usage behaviors in claims exploration
##\## Queries 2.00(2.48) 2.56(3.41) 10.00(0.25)
##\## Checkworthy slider changes 1.38(0.96) 6.44(5.84) 0.00(0.00)
##\## Query similarity slider changes 1.12(2.47) 1.44(2.63) 1.5(0.19)
##\## Selected claims 5.25(5.56) 5.44(3.52) 43.50(0.57)
##\## Final claims found checkworthy 1.44(1.03) 2.06(0.77) 18.0(0.04)
@@
##\## Queries 2.00(2.48) 1.06(1.81) 8.0(0.159) 2.56(3.41) 0.50(0.89) 0.0(0.01)
##\## Checkworthy slider changes 1.38(0.96) 1.31(1.25) 15.0(0.67) 6.44(5.84) 3.69(3.74) 23.0(0.12)
##\## Query similarity slider changes 1.12(2.47) 0.56(1.03) 6.0(0.68) 1.44(2.63) 0.38(0.72) 2.5(0.04)
##\## Selected claims 5.25(5.56) 4.06(2.43) 40.0(0.428) 5.44(3.52) 4.75(2.46) 48.0(0.78)
##\## Final claims found checkworthy 1.44(1.03) 1.88(1.02) 33.0(0.26) 2.06(0.77) 1.75(0.19) 33.5(0.20)
@@ 5.1.2. Design implications for personalized and efficient claim triage
First, when news events emerge and fact-checkers are uncertain about what topics or dis/misinformation narratives would be the central of fact-checking, they create general faceted filters (i.e., abstract prompts) to understand the semantics or potential harmful outcomes of problematic information. These criteria are frequently-used filters across different news events, as P14 explained “[they are] pretty broad category often used to filter out [less important] claims.” As fact-checkers become more familiar with specific dis/misinformation narratives, direct LLM-integrated search would be more helpful as it can retrieve specific claims based on precise narratives (i.e., targeted prompts).
@@
As search queries and/or LLM prompts become more complex, it becomes more important to explain how search results relate to different portions of an input query/prompt (i.e., algorithmic transparency). In IR, this is typically the domain of query-biased summarization (Sarwar et al., 2021) or search result snippets that serve to explain how each result relates to a user query. With our simple LLM integration, P2 reflected that “the result might not be directly explainable based on each of the things I wrote down [in the prompt]. ” P7 similarly remarked, “it would be nice to tell me why things were being arranged and returned in a certain way. For example, if I had looked up the spike protein in the prompt, I would have seen claims highlighting the spike protein.” While users today are very familiar with standard search engines and reformulating their queries to improve their search results, LLMs and prompt engineering are simultaneously far less familiar while also being seemingly far more complex.
@@ 5.2.1. Claim prioritization as a shift in fact-checking focus
However, due to the surge of online misinformation, post-hoc fact-checking has evolved into a more sophisticated digital practice, focusing on combating a variety of problematic information on social media (e.g., rumors, hoaxes, and propaganda) (Westlund et al., 2022). This digital practice also involves a larger group of stakeholders (Juneja and Mitra, 2022), including organizations solely checking online claims (e.g., Snopes, Lead Stories) and tech companies (e.g., Meta, Meedan) that provide automated fact-checking tools to support this effort (Arnold, 2020; Das et al., 2023; Dierickx et al., 2023). Claim prioritization now has evolved into a more intricate, nuanced, and challenging task, requiring alignment with the diverse values upheld by various stakeholders (Sehat et al., 2024; Liu et al., 2024; Procter et al., 2023).
@@ 6. Conclusion
Acknowledgements
We thank the professional fact-checkers who participated in our study for their valuable work and for making our research possible. This research was supported in part by the Knight Foundation and by Good Systems [^11] http://goodsystems.utexas.edu/, a UT Austin Grand Challenge to develop responsible AI technologies. The statements made herein are solely the opinions of the authors and do not reflect the views of the sponsoring agencies.
@@ Appendix G Other important checkworthy dimensions
Susceptibility (Babaei et al., 2022) was also invoked as another tie-breaking dimension beyond our study’s scope. As Sehat et al. (2024) note, susceptibility can serve as another indicator of harmfulness. P4 explained “If a social media user believed the false claim, it could potentially result in a more harmful outcome.
Section 5.3 further discusses our use of only four dimensions as a study limitation.
Appendix H Additional results
Likely false - - 3.69 — 4.31 0.0(0.00) 3.69 — 4.25 10.50(0.03)
Likely harmful - - - - 4.31 — 4.25 14.00(0.50)
Likely false - - 0.73 — 0.62 27.00(0.20) 0.73 — 0.39 8.50(0.01)
Likely harmful - - - - 0.62 — 0.39 11.00(0.02)

### v3

@@ Abstract.
Given the volume of potentially false claims online, claim prioritization is essential in allocating limited human resources available for fact-checking. In this study, we perceive claim prioritization as an information retrieval (IR) task: just as multidimensional IR relevance, with many factors influencing which search results a user deems relevant, checkworthiness is also multi-faceted, subjective, and even personal, with many factors influencing how fact-checkers triage and select which claims to check. Our study investigates both the multidimensional nature of checkworthiness and effective tool support to assist fact-checkers in claim prioritization. Methodologically, we pursue Research through Design combined with mixed-method evaluation.
Specifically, we develop an AI-assisted claim prioritization prototype as a probe to explore how fact-checkers use multidimensional checkworthy factors to prioritize claims, simultaneously probing fact-checker needs and exploring the design space to meet those needs. With 16 professional fact-checkers participating in our study, we uncover a hierarchical prioritization strategy fact-checkers implicitly use, revealing an underexplored aspect of their workflow, with actionable design recommendations for improving claim triage across multidimensional checkworthiness and tailoring this process with LLM integration.
@@
$c$s: Human-centered computing Interactive systems and toolsccs: Human-centered computing Empirical studies in HCIccs: Human-centered computing User centered design
@@ 3.1. Frame Research Goals
Figure 1. Our lab-based RtD method integrated into a classic double diamond process (Design Council, 2015). We document our design challenges as the research goals, described in Section 3.1. Steps of Discover and Define, i.e., how we explore and finalize design choices, are documented in Section 3.2. The step of Develop, i.e., how we create and deploy the prototype, is documented in Section $3.$ The step of Deliver, i.e., how we finally evaluate the design to gather use-inspired insights, is documented in Section 3.4. The iterative refinement of RtD evaluation protocol is documented in Appendix E.
@@ 3.3.2. Technical implementation
We implemented our models on the Streamlit server and built the front-end [^8] Streamlit: https://streamlit.io/, AG-Grid: https://www.ag-grid.com/ using Python, AG-Grid JS, and CSS. We used OpenAI’s gpt-3.5-turbo as our LLM to create customized classifiers [^9] While more advanced GPT models now exist, the log probability needed for building the ranking function was only available from gpt-3.5-turbo at the time of our implementation.. Codes can be accessed here  [^10] Codes: https://github.com/JialingJia/claim_prioritization.
@@ 3.4.2. Baseline condition
ID Gender Region Position Years Organizational Context Language Fact-checked
@@ 4.1.1. Perceived importance
Table 4. Mean (M), standard deviation (SD), and median statistics over participant data for different measures (rows, Section 3.4.4) and checkworthy dimensions (columns). Self-assessment data (top 3 rows) come from 5-point Likert scale answers, while observational data (bottom 3 rows) is drawn from interaction logs for the different UI sliders. Bold results indicate at least two checkworthy dimensions showed statistically significant differences (Friedman test at p ¡ 0.05). We observe significant differences between checkworthy dimensions for “Ease of finding” and “Overall weight” measures. Post-hoc test results are presented in Appendix H.
@@ 4.3. Targeted versus Abstract Prompting for Defining New Checkworthy Dimensions
I’ve done a lot of reporting on this [VAERS] and I’m familiar with the language, that’s like a huge subsection of COVID-19 claims… prompt like this helps me locate them.” (P13)
@@ 4.4.1. Usage behaviors in claims exploration
#\# Queries 2.00(2.48) 2.56(3.41) 10.00(0.25)
#\# Checkworthy slider changes 1.38(0.96) 6.44(5.84) 0.00(0.00)
#\# Query similarity slider changes 1.12(2.47) 1.44(2.63) 1.5(0.19)
#\# Selected claims 5.25(5.56) 5.44(3.52) 43.50(0.57)
#\# Final claims found checkworthy 1.44(1.03) 2.06(0.77) 18.0(0.04)
@@
#\# Queries 2.00(2.48) 1.06(1.81) 8.0(0.159) 2.56(3.41) 0.50(0.89) 0.0(0.01)
#\# Checkworthy slider changes 1.38(0.96) 1.31(1.25) 15.0(0.67) 6.44(5.84) 3.69(3.74) 23.0(0.12)
#\# Query similarity slider changes 1.12(2.47) 0.56(1.03) 6.0(0.68) 1.44(2.63) 0.38(0.72) 2.5(0.04)
#\# Selected claims 5.25(5.56) 4.06(2.43) 40.0(0.428) 5.44(3.52) 4.75(2.46) 48.0(0.78)
#\# Final claims found checkworthy 1.44(1.03) 1.88(1.02) 33.0(0.26) 2.06(0.77) 1.75(0.19) 33.5(0.20)
@@ 5.1.2. Design implications for personalized and efficient claim triage
First, when news events emerge and fact-checkers are uncertain about what topics or dis/ misinformation narratives would be the central of fact-checking, they create general faceted filters (i.e., abstract prompts) to understand the semantics or potential harmful outcomes of problematic information. These criteria are frequently-used filters across different news events, as P14 explained “[they are] pretty broad category often used to filter out [less important] claims.” As fact-checkers become more familiar with specific narratives, direct LLM-integrated search would be more helpful as it can retrieve specific claims based on precise narratives (i.e., targeted prompts).
@@
As search queries and/or LLM prompts become more complex, it becomes more important to explain how search results relate to different portions of an input query/prompt (i.e., algorithmic transparency). In IR, this is typically the domain of query-biased summarization (Sarwar et al., 2021) or search result snippets that serve to explain how each result relates to a user query. With our simple LLM integration, P2 reflected that “the result might not be directly explainable based on each of the things I wrote down [in the prompt]. ” P7 similarly remarked, “it would be nice to tell me why things were being arranged and returned in a certain way. For example, if I had looked up the spike protein in the prompt, I would have seen claims highlighting the spike protein.” While users today are adept at using standard search engines and refining their queries, LLM and prompt engineering remain relatively unfamiliar and appear more complex.
@@ 5.2.1. Claim prioritization as a shift in fact-checking focus
However, due to the surge of online misinformation, post-hoc fact-checking has evolved into a more sophisticated digital practice, focusing on combating a variety of problematic information on social media (e.g., rumors, hoaxes, and propaganda) (Westlund et al., 2022). This digital practice also involves a larger group of stakeholders (Juneja and Mitra, 2022), including organizations solely checking online claims (e.g., Snopes, Lead Stories) and tech companies (e.g., Meta, Meedan) that provide automated fact-checking tools to support this effort (Arnold, 2020; Das et al., 2023; Dierickx et al., 2023). Claim prioritization has evolved into a more intricate and nuanced task, requiring alignment with the diverse values of multiple stakeholders (Sehat et al., 2024; Liu et al., 2024; Procter et al., 2023).
@@ 6. Conclusion
Acknowledgments
We thank the professional fact-checkers who participated in our study for both their valuable work and for making our research possible. This research was supported in part by the Micron Foundation, the Knight Foundation [^12] https://knightfoundation.org/articles/connective-democracy-a-new-approach-for-building-bridges-in-american-society/ and by Good Systems [^13] http://goodsystems.utexas.edu/, a UT Austin Grand Challenge to develop responsible AI technologies. The statements made herein are solely the opinions of the authors and do not reflect the views of the sponsoring agencies.
@@ Appendix G Other important checkworthy dimensions
Susceptibility (Babaei et al., 2022) was also invoked as another tie-breaking dimension beyond our study’s scope. As Sehat et al. (2024) note, susceptibility can serve as another indicator of harmfulness. P4 explained “If a social media user believed the false claim, it could potentially result in a more harmful outcome.” Section 5.3 further discusses our use of only four dimensions as a study limitation.
Appendix H Additional statistical results
@@
Likely false - - 3.69 — 4.31 0.0(0.00) 3.69 — 4.25 10.50(0.03)
Likely harmful - - - - 4.31 — 4.25 14.00(0.50)
Likely false - - 0.73 — 0.62 27.00(0.20) 0.73 — 0.39 8.50(0.01)
Likely harmful - - - - 0.62 — 0.39 11.00(0.02)